OpenStudy (anonymous):

Use parametric equations to model the path of a rider on the wheel. You want to end up with parametric equations, using t, that will give you the position of the rider every minute of the ride. Graph your results on a graphing calculator. An observation wheel has a diameter of 120 meters and sits 16 meters above the ground. It rotates in a counterclockwise direction, making one complete revolution every 32 minutes. Place your coordinate system so that the origin of the coordinate system is on the ground 110 meters to the left of the wheel.

OpenStudy (loser66):

Where are you stuck?

OpenStudy (loser66):



OpenStudy (loser66):

Consider the x-coordinate. Let A is the beginning point when the wheel moves. The wheel, is originally far from the origin 110m, so that x (t) = 110 +.....something

OpenStudy (anonymous):

Im not sure where to start with the problem

OpenStudy (loser66):

that something is calculated by : The diameter of the wheel is 120, hence the perimeter is 120pi

OpenStudy (anonymous):

I thought it was x = 110+ something but I dont know how to find the rest of the problem

OpenStudy (loser66):

And it takes 32 minutes to finish 1 revolution, that shows at the time t, it goes \(\dfrac{120\pi}{32}t\), So, to x, it is \(x = 110 + \dfrac{120\pi}{32}t\)

OpenStudy (anonymous):

oh okay.. so do I get the 120pi and divide by something?

OpenStudy (anonymous):

sorry, I didnt see your post before sending my question

OpenStudy (loser66):

Do the same with y, and you get both parametric equation.

OpenStudy (loser66):

It's ok, as long as you get it, nothing is important. :)